# Bridging Today and the Future of Humanity: AI Safety in 2024 and Beyond

### v2

@@ Abstract
The growing prevalence of generative AI inevitably raises concerns regarding the associated risks and safety implications, which catalyzes significant progress in AI safety. However, as this field thrives, a critical question emerges: Are our current efforts aligned with the broader perspective of human history and civilization? This paper presents a blueprint for an advanced human society and leverages this vision to guide contemporary AI safety efforts. It outlines a future where the Internet of Everything becomes reality, and create a roadmap of significant technological advancements towards this envisioned future. For each stage of the advancements, this paper forecasts potential AI safety issues that humanity may face. By projecting current efforts against this blueprint, we examine the alignment between the present efforts and the long-term needs. This paper identifies gaps in current approaches and highlights unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s, addressing critical areas that must not be overlooked in shaping a responsible future for AI development. This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby fostering AI’s role in promoting a more secure and sustainable future for human civilization.
@@ 1 Introduction
The rapid development of AI and Large Language Models (LLMs) has fostered extensive progress in AI safety, specifically, LLM safety [171, 143, 51, 230, 90, 37]. Researchers have been dedicated to addressing potential security and privacy risks in LLM lifecycle, aiming to align AI behaviors with human values and prevent misuse of LLMs, inappropriate outputs, information leakage, etc. However, despite the intense focus on LLM safety, a critical question emerges: Are current AI safety efforts truly aligned with the long-term evolution of human civilization, or are they simply addressing the concerns of today?
@@
One fundamental reason for this uncertainty lies in the probabilistic nature of current AI [146]. Despite their impressive capabilities in natural language processing and problem-solving [270, 265, 54, 38], today’s AI, including advanced LLMs [286, 260, 231], falls short of what could be considered genuine intelligence. These systems heavily rely on vast training datasets to function effectively, and lack true understanding, consciousness, self-awareness, and the ability to reason in ways comparable to human cognition. They are, in essence, highly sophisticated pattern recognition and prediction machines, rather than entities capable of abstract thought and genuine reasoning. Moreover, recent studies have raised concerns about the reasoning abilities of LLMs, which may actually be a form of approximate retrieval from their training data and the deductive closure of the training data [96, 239]. While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it fundamentally differs from human reasoning that involves abstract thinking, causal understanding, and the ability to generalize from limited examples.
@@
The unresolved issue of sustainable energy is another critical factor that challenges our current efforts to AI safety [5]. Current AI fails to fully represent a third industrial revolution from the perspective of human history, as the fundamental energy issue remains unsolved. Historically, the two industrial revolutions that have profoundly shaped human civilization were driven by revolutionary energy innovations, with the First Industrial Revolution fueled by steam and coal, and the Second Industrial Revolution characterized by a new wave of technological innovations powered by electricity and petroleum [76]. However, current AI, rather than solving energy issues, consumes a significant amount of energy. Training GPT-4 consumed over 50,000 MWh, 10,353.5 tons of CO2 equivalent and approximately 0.02% of California’s annual electricity generation [7, 39, 169]. Also, inference with LLMs is computationally intense, e.g., a single query to GPT-4 consumes 0.001 to 0.01 kWh, approximately 15x energy than a Google query [140, 72]. Given that ChatGPT alone has over 200 million weekly active users and receives over 1.54 billion page visits monthly [213], when scaled to billions of queries, the energy consumption becomes substantial. The significant energy requirements of current AI systems raise serious sustainability concerns and limiting their deployment throughout the human society. Until we overcome the energy constraints, AI’s ability to reshape society remains restricted.
@@
These concerns suggest that our efforts on AI safety today may fail to capture deeper insights that could emerge as AI evolves, potentially neglecting technological challenges that ultimately shape the impact of AI on human civilization. This raises several critical questions:
  • Are we focusing on the right problems that align with the long-term AI development?
@@
  • Are we on the right track for addressing real challenges in AI safety?
“The Moon and Sixpence.” This paper envisions an ultimate advanced society in a distant future of human civilization, and anticipates potential technological advancements to guide today’s AI safety research. In this advanced future, energy issues have been resolved, and intelligent chips and brain-computer interfaces have been fully developed, which enables advanced intelligent robots to integrate into everyday life. Breakthroughs in foundational theories, such as Einstein’s relativity [57, 58] and Quantum Mechanism [59, 279], have catalyzed revolutionary progress in AI, and humans and robots are connected in an interconnected network, i.e., the Internet of Everything; see Figure 1. While this vision may be beyond the horizon of several decades, by reviewing the present landscape of AI safety against this far future, we can gain deeper insights into the potential and limitations of our current approaches, as well as specific missions for AI safety researchers in the 2020s. It reminds us that, while addressing today’s immediate concerns, we should avoid being constrained by these short-term solutions and look toward the future to guide our present efforts.
Takeaways: i) The term “AI safety” can be extended to “AI quality assurance” to encompasses more comprehensive concepts and possibilities in the fast evolving AI landscape. ii) Deeper collaboration with experts from diverse fields is essential to align AI with human values better. iii) AI workflow customization and user-AI interactions are important, especially in dynamic contexts and across different user groups.
@@ 2 Related Works
Multiple AI safety surveys and benchmarks have been published in recent years. These studies focus on different aspects of AI safety, such as jailbreaking, red teaming, guardrail, ethics, etc. A summary of these works is presented in Table 1, which provides an overview of the related works and their focus within the field of AI safety.
Among these works, [124] provides comprehensive insights into building trustworthy AI systems by focusing on reliability, safety, fairness, robustness, and adherence to social norms. [37] addresses AI safety concerns from the perspective of different entities, e.g., data and prompts, in LLM lifecycle, and discusses data safety, model safety, prompt safety, alignment, and the complexities of scaling.  [102] provides valuable enterprise insights into the practical challenges and opportunities for AI safety in industries. [67] provides a systematic analysis of the ethical risks of advanced AI assistants, and discusses their potential influence on users, the risks of misuse, strategies for enhancing human-AI interactions, and the broader implications for human society. [43] proposes the concept of “Guaranteed Safe” (GS) AI, emphasizing the need for formal, verifiable, and auditable safety guarantees to ensure robust AI behavior.
[11] proposes a technical roadmap towards next-generation human-centered AI safety that aims to anticipate and prevent potential hazards in the interactions between AI systems and users. Specifically, the paper bridges the fields of AI safety and control systems, and highlights the open challenges at their intersection. The authors emphasize the importance of leveraging control-theoretic safety methodologies alongside the rich interaction models of generative AI, suggesting that a deeper understanding of the feedback loop between AI outputs and human behavior is essential for achieving meaningful safety assurances.
Different from existing works, this paper envisions an advanced society in the future of human civilization, and use this vision to guide current AI safety research. By comparing the current landscape of AI safety with this distant future, this work identifies both the potential and limitations of today’s approaches and suggests specific missions for AI safety researchers in the 2020s. This perspective encourages maintaining a forward-looking vision while resolving immediate concerns, such that we can ensure that today’s solutions are aligned with the long-term evolution of AI and human society.
@@ 3 From the New World
In this future, humanity has entered an unprecedented era. Scientists have achieved breakthroughs in energy production with technological innovations. The resulting flood of abundant, clean energy has revolutionized human civilization beyond the our ancestors’ wildest dreams, reshaping the world and initiating a new chapter of human civilization.
@@ 3.1 Imagine the Future: A Blueprint
The transformative energy revolution has catalyzed human civilization from every aspect. Foundational theories of science have undergone revolutionary breakthroughs, with major advancements in nanotechnology [136, 19, 61], Quantum Mechanics [59, 279], and Einstein’s relativity theories [57, 58]. The world has been reshaped by innovative technologies such as intelligent chips [86, 115], brain-computer interfaces [147, 81, 254], holographic technology [22, 273, 15], and advanced 3D printing [205, 73], bringing changes in daily life of miraculous.
@@
Transportation has evolved dramatically. While simple mechanical vehicles, such as fixed-track trains and high-speed rail [144, 179, 109], still exist due to their efficient and straightforward design, they operate at astonishing speeds with high stability, powered by efficient and limitless energy. Advanced autonomous vehicles and aircraft are widespread and are controlled by intelligent chips, eliminating the need for any human intervention. Also, due to breakthroughs in Einstein’s relativity theories [57, 58] and Quantum Mechanics [59, 279], humanity has unlocked the potential of instant transportation that allow for long-distance travel in short time that transcends the imagination of the 21st century.
@@
Advanced brain-computer interfaces [147, 81, 254] and intelligent microchips [86, 115] has revolutionized the way people perceive, which expands the boundaries of human cognition. No longer are people operating devices and asking questions in search engines or LLMs; the embedded intelligence chips help them to gather more comprehensive information and interact with the surrounding environment. Meanwhile, communication over long distances has transcended the limitations of cables and the outdated internet; information, thoughts, emotions, and sensations flow through a new medium due to the advancement in quantum computing [77, 198].
@@
Intelligent robots have been integrated into every aspect of modern society and play important roles of daily life; see Figure 1. Different from their “data-hungry machine intelligent ancestors” that relies on huge amount of training data and statistical patterns in the early 21st century [131, 286], these intelligent robots perceive and learn through intelligent chips, which are real artificial intelligence and possess genuine understanding and adaptability, much like a human child discovering the world for the first time. Also, their functionalities are customizable through specialized chips based on user needs. Robots can engage in complex interactions with people and adapt their capabilities and expertise to suit their human counterparts’ requirements.
@@
In this brave new world, humans and robots exist within a vast, interconnected perceptual network, where the boundaries between the digital and physical worlds have been blurred, i.e., Internet of Everything. Such a network integrates objects with embedded intelligent chips, such as smart devices, autonomous systems, intelligent robots, and humans cognition, which allows for unprecedented levels of efficiency, personalization, and adaptability and fundamentally reshapes the way people interact with the world.
@@ 3.2 Retrospective on AI Safety in the New World
In a world where the Internet of Everything has become a reality, intelligent robots are customized to meet diverse user needs and seamlessly integrate into daily life; see Figure 1. Consequently, the concept of AI safety has been redefined to focus more on the quality of services they provide, i.e., emphasizing comprehensive quality assurance. Unlike today’s command-and-response interactions between humans and LLMs, future robots can engage in sophisticated reasoning, emotional cognition, and possess appropriate levels of autonomy. They will also interact dynamically with other entities, including humans and fellow robots, within the interconnected system.
@@
The Internet of Everything scenarios necessitate the robots to handle complex interactions, demanding a sophisticated understanding of the environment for decision-making. Care robots, for instance, should not only assist physically but also recognize emotions and offer emotional support or suggestions, while cleaning robots need to assess whether a room requires more work based on its cleanliness level. Such advanced interactions require deep understanding of the environments, demanding real-time information retrieval and processing, which is far beyond what can be achieved with simple code logic. Robots must continuously collect and analyze data from their environment, and integrate external inputs with contextual knowledge to facilitate real-time decision-making, such that they can deliver high-quality services that align precisely with human needs and expectations.
@@
Customization is a key aspect of future robots for providing personalized services to meet individuals’ requirements. Even robots of the same type should be capable of adapting their services to different individuals. Teaching robots should adjust educational materials and teaching methods to capitalize on each student’s unique talents and strengths, while cleaning robots should intelligently schedule their works to minimize disruption to other humans in the household, e.g., avoiding occupied workspace. Such customization aligns the robots’ capabilities with specific user preferences, and ensures that the robots serve as effective assistants across a wide range of daily tasks and environments.
@@
Finally, the effectiveness of intelligent robots requires a balance between robot autonomy with human oversights. Too little autonomy would burden users with continuous supervision, while too much autonomy could raise safety concerns or result in unintended behaviors. For instance, a care robot may operate with high autonomy when monitoring vital signs or providing routine care, but would require human intervention for critical decisions. Conversely, while the tasks of cleaning robots are generally routine, they must still adapt to specific circumstances, such as recognizing when an area is occupied by people and postponing cleaning activities. An appropriate level of robot autonomy optimizes the division of labor between humans and robots, ensuring that robots function effectively and safely while enhancing our daily lives without compromising human control over these intelligent robots.
@@ 4 Bridging Today and the Future
This section presents an ambitious roadmap towards the aspirations for the advanced future, exploring key milestones in the evolution of AI and human civilization. By considering potential paradigm shifts in fundamental theories and breakthroughs in technology, we feature this fascinating journey into three eras, i.e., i) the data-driven era where AI learns knowledge from vast training datasets, ii) the neural augmentation era where human-machine interface [147, 81, 254] and intelligent chips [86, 115] extend the boundary of human cognition, and iii) the revolutionary era with a breakthrough in energy production that catalyzes transformative changes across every aspect of human society; see §4.1, §4.2, and §4.3, respectively. We then examine the evolving paradigms of AI safety throughout this journey, analyzing the potential focuses and challenges of each era and exploring the impacts of technological advancement on AI safety in §$4.$

### v3

@@ Abstract
The advancements in generative AI inevitably raise concerns about the associated risks and safety implications, which, in return, catalyzes significant progress in AI safety. However, as this field continues to evolve, a critical question arises: are our current efforts aligned with the long-term goal of human history and civilization? This paper presents a blueprint for an advanced human society and leverages this vision to guide contemporary AI safety efforts. It outlines a future where the Internet of Everything becomes reality, and creates a roadmap of significant technological advancements towards this envisioned future. For each stage of the advancements, this paper forecasts potential AI safety issues that humanity may face. By projecting current efforts against this blueprint, we examine the alignment between the present efforts and the long-term needs. We also identify gaps in current approaches and highlight unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s, addressing critical areas that must not be overlooked in shaping a responsible and promising future of AI. This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby promoting a more secure and sustainable future in human civilization.
@@ 1 Introduction
The rapid developments of AI and Large Language Models (LLMs) have fostered extensive progress in AI safety, specifically, LLM safety [175, 147, 54, 235, 93, 39]. Researchers have been dedicated to addressing potential security and privacy risks in AI lifecycle, aiming to align AI behaviors with human values and prevent the misuse of AI models, inappropriate outputs, information leakage, etc. However, despite the intense focus on AI safety, a critical question emerges: are current AI safety efforts truly aligned with the long-term evolution of human civilization, or are they simply addressing the immediate concerns of the 2020s?
@@
One fundamental reason for this uncertainty lies in the probabilistic nature of current AI [150]. Despite their impressive capabilities in natural language processing and problem-solving [280, 274, 57, 40], today’s AI, including advanced LLMs [297, 268, 236], falls short of what could be considered as “genuine intelligence”. Current AI models rely heavily on vast training datasets to function effectively, yet lack true understanding, consciousness, self-awareness, and reasoning abilities comparable to human cognition. They are, in essence, highly sophisticated pattern recognition and prediction machines, rather than entities with authentic logical capabilities. Recent studies have raised concerns about the reasoning abilities of LLMs, which may be a form of approximate retrieval from their training data and the deductive closure of the training data [100, 245]. While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it differs fundamentally from human reasoning that involves abstract thinking, causal understanding, and generalizing from limited examples.
@@
The energy issue is another critical factor that challenges our current efforts on AI safety [5]. Current AI fails to fully represent a third industrial revolution from the perspective of human history, as the fundamental energy issue remains unsolved. Historically, the two industrial revolutions that have profoundly shaped human civilization were driven by revolutionary energy innovations, with the First Industrial Revolution fueled by steam and coal, and the Second Industrial Revolution characterized by a new wave of technological innovations powered by electricity and petroleum [79]. However, current AI, rather than solving energy issues, consumes a significant amount of energy. Training GPT-4 consumed over 50,000 MWh, 10,353.5 tons of CO2 equivalent, and approximately 0.02% of California’s annual electricity generation [7, 41, 173]. Also, inference with LLMs is computationally intense, e.g., a single query to GPT-4 consumes 0.001 to 0.01 kWh, approximately 15x energy than a Google query [144, 75]. Given that ChatGPT alone has over 200 million weekly active users and receives over 1.54 billion page visits monthly [218], when scaled to billions of queries, the energy consumption becomes substantial. The significant energy consumption of current AI systems raises concerns about their sustainability, limiting their widespread deployment across human society. Until humanity overcomes the energy constraints, the potential of AI to reshape society will remain limited and uncertain.
@@
These concerns indicate that our current efforts in AI safety might overlook more complex issues that could arise as AI continues to evolve. We potentially fail to capture deeper insights that will ultimately shape the impact of AI on human civilization. This raises several critical questions:
  • Are we focusing on the right problems that align with the long-term advancements of AI?
@@
  • Are we on the right track of addressing real challenges in AI safety?
“The Moon and Sixpence.” This paper envisions an ultimate advanced society in a distant future of human civilization, and anticipates potential technological advancements to guide today’s AI safety research. In this advanced future, energy issues have been resolved, and breakthroughs in foundational theories, such as Einstein’s relativity [60, 61] and Quantum Mechanism [62, 289], have catalyzed revolutionary progress in AI. Intelligent chips [89, 119] and brain-computer interfaces [84, 151, 262] have been fully developed, and advanced intelligent robots are integrated into everyday life. Humans and robots are connected in an interconnected network, i.e., the Internet of Everything; see Figure 1. While this vision may be beyond the horizon of several decades, by reviewing the present landscape of AI safety against this far future, we can gain deeper insights into the potential and limitations of our current efforts, as well as specific missions for AI safety researchers in the 2020s. It reminds us that, while addressing today’s immediate concerns, we should avoid being limited by short-term solutions, and instead, look toward the future to guide our present efforts.
Takeaways: i) The term “AI safety” can be extended to “AI quality assurance” to encompasses more comprehensive concepts and possibilities in the fast evolving AI landscape. ii) Deeper collaborations between experts from different fields are essential to align AI with human values better. iii) AI and AI safety workflow customizations are crucial to ensure that AI systems can adapt to diverse user requirements while maintaining safety and reliability across different use cases. iv) Exploring and understanding interactions between users and AI are crucial for enhancing user experiences, especially in dynamic contexts across different user groups.
@@ 2 Related Works
Multiple AI safety surveys and benchmarks have been published in recent years [106, 8, 126, 39, 189, 128, 85, 282]. These studies focus on different aspects of AI safety, such as jailbreaking, red teaming, guardrail, ethics, etc [128, 282, 146]. A summary of these works is presented in Table 1, which provides an overview of the related works and their focus within the field of AI safety.
Among these works, [128] provides comprehensive insights into building trustworthy AI systems by focusing on reliability, safety, fairness, robustness, and adherence to social norms. [39] addresses AI safety concerns from the perspective of different entities, e.g., data and prompts, in LLM lifecycle, and discusses data safety, model safety, prompt safety, alignment, and the complexities of scaling.  [106] provides valuable enterprise insights into the practical challenges and opportunities for AI safety in industries. [70] provides a systematic analysis of the ethical risks of advanced AI assistants, and discusses their potential influence on users, risks of misuse, strategies for enhancing human-AI interactions, and the broader implications for human society. [45] proposes the concept of “Guaranteed Safe” AI, emphasizing the need for formal, verifiable, and auditable safety guarantees to ensure robust AI behavior. [11] proposes a technical roadmap towards next-generation human-centered AI safety that aims to anticipate and prevent potential hazards in the interactions between AI systems and users. This paper bridges the fields of AI safety and control systems, and highlights the open challenges at their intersection. The authors emphasize the importance of leveraging control-theoretic safety methodologies alongside the rich interaction models of generative AI, suggesting that a deeper understanding of the feedback loop between AI outputs and human behavior is essential for achieving meaningful safety assurances.
Different from existing works, this paper envisions an advanced society in the future of human civilization, and use this vision to guide current AI safety efforts. By aligning the current landscape of AI safety with this distant future, we identify the potential and limitations of today’s efforts, and suggest specific missions of AI safety researchers in the 2020s. This paper encourages maintaining a forward-looking vision while resolving immediate concerns, such that we can ensure that today’s solutions are aligned with the long-term evolution of AI and human society.
@@ 3 From the New World
In this future, humanity has entered an unprecedented era, driven by technological breakthroughs in energy generation. The flood of clean energy has revolutionized human society beyond the wildest dreams of our ancestors, reshaping the world and initiating a new chapter of human history.
@@ 3.1 Imagine the Future: A Blueprint
The energy revolution has transformed nearly every aspect of human society. Foundational theories have experienced groundbreaking advancements, particularly in fields such as nanotechnology [140, 19, 64], quantum mechanics [62, 289], and Einstein’s relativity theories [60, 61]. Innovative technologies, such as intelligent chips [89, 119], brain-computer interfaces [151, 84, 262], holographic technology [22, 283, 15], and advanced 3D printing [210, 76], have reshaped the world, bringing about profound changes in daily life that were once beyond our imagination.
@@
Transportation has evolved dramatically. While simple mechanical vehicles, such as fixed-track trains and high-speed rails [148, 183, 113], still exist due to their efficient and straightforward design, they operate at astonishing speeds with high stability, powered by abundant efficient energy. Advanced autonomous vehicles and aircraft, controlled by intelligent chips, are widespread, eliminating the need for human intervention. Also, with breakthroughs in Einstein’s relativity theories [60, 61] and quantum mechanics [62, 289], humanity has unlocked the potential of instant transportation that allow for long-distance travel in short time that transcends the imagination of the 21st century.
@@
Advancements in brain-computer interfaces [84, 151, 262] and intelligent microchips [89, 119] have revolutionized the way people perceive and expanded the boundaries of human cognition. No longer are people operating devices and asking questions in search engines or AI models such as LLMs; the embedded intelligence chips help people to gather comprehensive information and interact with the surroundings. Also, communication over long distances has transcended the limitations of the outdated cables and the internet; information, thoughts, emotions, and sensations flow through a new medium due to the advancements in quantum computing [80, 203].
@@
Intelligent robots have been integrated deeply into human society and play important roles in daily life; see Figure 1. Different from their “data-hungry machine intelligent ancestors” that rely on huge amount of training data and statistical patterns in the early 21st century [135, 297], these intelligent robots perceive and learn through intelligent chips. They are real artificial intelligence that possess genuine abilities of learning, understanding, and adapting, much like a human child discovering the world for the first time. Their functionalities are customizable through specialized chips based on user needs, and they can engage in complex interactions with people and adapt their capabilities and expertise according to their human counterparts’ requirements.
@@
In this brave new world, humans and robots exist within a vast, interconnected perceptual network, where the boundaries between the digital and physical worlds are blurred, i.e., the Internet of Everything. Such networks integrate objects embedded with intelligent chips, e.g., smart devices, autonomous systems, intelligent robots, and humans, allowing for unprecedented levels of efficiency and adaptability and fundamentally transforming the way people interact with the world.
@@ 3.2 Retrospective on AI Safety in the New World
In a world where the Internet of Everything has become a reality, intelligent robots are customized to meet diverse user needs and are seamlessly integrated into daily life; see Figure 1. Consequently, the concept of AI safety has been redefined to prioritize the quality of services provided by the intelligent robots, i.e., emphasizing comprehensive quality assurance rather than solely focusing on safety, security, and privacy risks. Unlike today’s command-and-response interactions between humans and AI models, future robots can engage in sophisticated reasoning and emotional cognition, and possess appropriate levels of autonomy. They will also interact dynamically with other entities, including humans and fellow robots, within the interconnected network.
@@
The Internet of Everything scenarios necessitate the robots to handle complex interactions, thus requiring a sophisticated understanding of the surroundings for decision-making. Care robots, for instance, should not only assist physically but also recognize emotions and offer emotional support or suggestions, while cleaning robots need to assess whether a room requires more work based on its cleanliness level. Such advanced interactions requires the intelligent robots to retrieve and process information in real-time, which is far beyond what can be achieved with simple code logic. Robots must continuously collect and analyze data from the surroundings and integrate external inputs with contextual knowledge to facilitate real-time decision-making, such that they can deliver high-quality services that align with human expectations.
@@
Customization is a key aspect of future robots for providing personalized services to meet individuals’ requirements. Even robots of the same type should be able to adapt their services to different individuals. Teaching robots should adapt educational materials and teaching methods to leverage each student’s unique talents and strengths, while cleaning robots should intelligently schedule their tasks to minimize disruption to household members, such as avoiding occupied workspace. Such customization aligns the robots’ capabilities with specific user needs, ensuring that the robots serve as effective assistants across a wide range of daily tasks and complex environments.
@@
Finally, the effectiveness of intelligent robots requires a balance between robot autonomy with human oversights. Too little autonomy would burden users with continuous supervision, while too much autonomy could raise safety concerns or result in unintended behaviors. A care robot may operate with high autonomy when monitoring vital signs or providing routine care, but would require human intervention for critical decisions. Conversely, while the tasks for cleaning robots are typically routine, they still need to adapt to specific situations, such as detecting when an area is occupied and postponing cleaning activities accordingly. An appropriate level of robot autonomy ensures an effective division of labor between humans and robots, allowing robots to operate efficiently and safely while enhancing our daily life without compromising human control over these intelligences.
@@ 4 Bridging Today and the Future
This section outlines an ambitious roadmap to the advanced future, highlighting key milestones in the evolution of AI and human civilization. By exploring potential paradigm shifts in fundamental theories and technological breakthroughs, this section categorizes this journey into three eras, i.e., i) the data-driven era, in which AI learns from vast training datasets, ii) the neural augmentation era, where human-machine interface [151, 84, 262] and intelligent chips [89, 119] expand the boundaries of human cognition, and iii) the revolutionary era, characterized by a breakthrough in energy production that catalyzes transformative changes across all aspects of human society; see §4.1, §4.2, and §4.3, respectively. We then examine the evolving paradigms of AI safety throughout this journey, analyzing the focuses and challenges of each era while exploring the impacts of technological advancement on AI safety in §4.4.